If the simple pendulum is brought from one place to another, its time period gets increased. How
1
Expert's answer
2020-03-11T14:21:08-0400

Simple pendulum has a period that depends on the acceleration due to gravity:


"T=2\pi\sqrt{\frac{l}{g}}."

We see that the period can increase if g decreases. Where can it decrease? At the equator because the earth's gravitational attraction on the bob is affected by the highest speed of rotation at the equator which decreases the value of g to 9.78 m/s2 compared to 9.832 m/s2 at the poles.

Also, g can be decreased at big heights above the sea level (again according to the law of universal gravitation )